OpenStudy (anonymous):

The point (0, -3.5) lies on the line 11x – 6y = 21. Determine if this point is a solution of the system 11x – 6y = 21 9x + 8y = 43

OpenStudy (anonymous):

substitude the values it satisfies the equation so it is a solution

OpenStudy (anonymous):

u want to clear explanation?

OpenStudy (anonymous):

. means multiplcation in notation 11.0-(-) 6.7/2=0-(-)21=21

OpenStudy (anonymous):

x=0 y=7/2 substitude